# Why Can’t I Open My Drawer? Mitigating Object-Driven Shortcuts in Zero-Shot Compositional Action Recognition

### v1

@@ D.1 Comprehensive ablation results
(b) Temporal/Static splits from Sevilla et al. [31]
@@
Figure 10: Performances on Temporal/Static split of Sth-com. We evaluate the models on Sth-com [16] using both (a) our reconstructed splits and (b) the splits from Sevilla et al [31]. We utilize both original and temporally shuffled inputs to assess the model’s temporal modeling capability and its reliance on static cues. A larger performance gap between original and shuffled inputs indicates that the model predicts verbs more based on temporal dynamics rather than on static cues.
Effects of hyperparameters of RCORE.
In Table 9 (f), we demonstrate that applying augmentation with a probability of 0.5 within a batch is the most effective, achieving a 0.4 point improvement in H.M. over the baseline. In Table 9 (g), we observe that introducing the explicit temporal order modeling task via TORC starting from 5 epoch, after initial training, is effective. In particular, the strategy of linearly increasing the loss weight from epoch 5 to 10 is the most effective, resulting in 1.7 point boost in H.M. compared to the baseline. Finally, in Table 9 (h), we analyze the impact of scaling the weight of our Margin loss, revealing a trade-off between improved unseen composition accuracy and decreased seen composition accuracy. We observe that applying Margin loss starting from epoch 15, the point at which the model begins to overfit the training data as shown in Figure 3 of the main paper, is more effective than starting from the earlier epoch. We also find that linearly scaling the loss weight over 5 epochs is the most effective strategy to maintain the model’s performance.
D.2 Analysis of confidence distribution in composition classification
(a) Sth-com results
(b) EK100-com results
Figure 11: Conditional modeling overfits to co-occurrence statistics. We track how the composition prediction confidence of C2C [16] evolves during training both on Sth-com [16] and EK100-com datasets. As training progresses, C2C [16] increasingly ignores input evidence and misclassifies unseen compositions as seen ones.
In Figure 11, we track how the composition prediction confidence of C2C [16] evolves during training both on Sth-com [16] and EK100-com datasets. Using the test set of both datasets, we compare two cases of the model’s confidence values: Seen $\to$ Seen and Unseen $\to$ Seen. Given the seen/unseen composition set $\mathbb{Y}_{\text{seen}}$$\mathbb{Y}_{\text{unseen}}$ and the ground-truth/model-predicted composition label $\mathbf{y}$$\hat{\mathbf{y}}$, Seen $\to$ Seen refers to the case where $\mathbf{y}_{i}\in\mathbb{Y}_{\text{seen}}\land\hat{\mathbf{y}}_{i}\in\mathbb{Y}_{\text{seen}}$ and Unseen $\to$ Seen refers to the case where $\mathbf{y}_{i}\in\mathbb{Y}_{\text{unseen}}\land\hat{\mathbf{y}}_{i}\in\mathbb{Y}_{\text{seen}}$. Across both datasets, we observe that the model increasingly classifies unseen compositions as seen compositions as training progresses. The observation suggests that, without careful design, models may be prone to ignore input cues and instead over-rely on verb–object co-occurrence statistics present in the training data.
D.3 Temporal/Static split of Sth-com
Table 10: Ablation study on the EK100-com. We provide the results of the ablation study on EK100-com presented in the main paper. In every experiment, we use C2C [16] with CLIP-B/16 [28] as our baseline. We report performance on both seen and unseen compositions, along with their harmonic mean (H.M.).
(a) Effects of VOCAMix and TORC. VOCAMix TORC Verb Object Composition (OW) Seen Unseen Seen Unseen Seen Unseen H.M. 66.19 49.71 56.61 47.48 42.72 22.38 29.38 65.55 52.15 57.09 46.87 43.01 23.81 30.65 65.90 52.31 57.28 46.74 43.48 23.67 30.65 66.21 53.51 56.97 47.52 43.53 25.02 31.78
(b) Effects of each TORC loss term. $L_{\cos}$ $L_{\text{ent}}$ Verb Object Composition (OW) Seen Unseen Seen Unseen Seen Unseen H.M. 66.19 49.71 56.61 47.48 42.72 22.38 29.38 65.43 52.11 56.54 47.44 42.67 24.24 30.92 66.00 51.06 56.64 47.37 42.82 22.53 29.53 65.90 52.31 57.28 46.74 43.48 23.67 30.65
@@ D.3 Temporal/Static split of Sth-com
We present extended results corresponding to the experiment shown in Figure 5 (c) of the main paper, including all versions of the Temporal/Static split. Inspired by prior work [31, 39], we rigorously evaluate the temporal modeling capabilities of both the baseline (C2C [16]) and our RCORE by partitioning the Sth-com [16] dataset into Temporal and Static splits. Sevilla et al. [31] originally defined these splits based on cognitive science experiments on Something-Something V2 [12] verb classes: a Temporal split (18 verbs) where temporal information matters, and a Static split (18 verbs) where such information is redundant. However, noting that this Temporal split includes verbs requiring complex temporal reasoning such as ‘Folding’, we reconfigure the splits based on verb semantics. Specifically, we categorized verbs based on this criterion: those requiring temporal information for discrimination are assigned to the Temporal split (62 verbs), while verbs distinguishable by momentary keyframes are assigned to the Static split (80 verbs). The indices of the verbs in the Temporal split are as follows: 0, 1, 5, 6, 14, 25, 26, 29, 30, 32, 33, 34, 35, 38, 39, 40, 41, 42, 43, 44, 45, 46, 58, 59, 60, 61, 65, 66, 67, 68, 72, 73, 74, 75, 76, 78, 79, 85, 86, 90, 91, 92, 98, 100, 103, 109, 122, 123, 131, 134, 136, 137, 139, 140, 145, 146, 147, 148, 149, 150, 153, 154. The indices of the verbs in the Static split are as follows: 2, 3, 8, 9, 10, 11, 12, 13, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 27, 28, 31, 36, 37, 47, 48, 49, 50, 51, 52, 53, 54, 55, 77, 80, 81, 87, 88, 93, 94, 95, 96, 97, 99, 101, 102, 104, 105, 106, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 126, 127, 128, 129, 130, 132, 133, 135, 138, 141, 142, 143, 151, 152, 155, 156, 157, 158, 159, 160. To balance the sample sizes between the Temporal and Static splits, we excluded a few verbs with the fewest samples from the Temporal split. We then utilize both original and temporally shuffled inputs to assess the model’s temporal modeling capability and its reliance on static cues. In Figure 10, we present the results for the Temporal/Static splits across seen/unseen compositions. We also provide results applying the original Temporal/Static verb split [31] to Sth-com [16]. In all cases, RCORE exhibits lower performance on shuffled inputs but outperforms the baseline on original inputs. This indicates that RCORE predicts verbs without relying on static cues, leading to robust improvements in verb performance regardless of the split.
D.4 Ablation studies on EK100-com
In Table 10, we provide the ablation study results conducted on the EK100-com dataset, complementing the results discussed with the Sth-com [16] dataset in the main paper. In Table 10 (a), we demonstrate that our framework, integrating VOCAMix for sparsity resolution and TORC for explicit temporal modeling, outperforms the baseline on EK100-com, achieving 3.8 points higher verb accuracy in unseen compositions than the baseline. In Table 10 (b), using only $L_{cos}$ in TORC yields the highest harmonic mean of composition accuracy on EK100-com.

### v2

@@ 0.E.4 Results with validation-set-tuned bias calibration
(b) Temporal/Static splits from Sevilla et al. [laura2021temporal]
@@
Figure 14: Performances on Temporal/Static split of Sth-com. We evaluate the models on Sth-com [li2024c2c] using both (a) our reconstructed splits and (b) the splits from Sevilla et al [laura2021temporal]. We utilize both original and temporally shuffled inputs to assess the model’s temporal modeling capability and its reliance on static cues. A larger performance gap between original and shuffled inputs indicates that the model predicts verbs more based on temporal dynamics rather than on static cues.
0.E.5 Results with large video-pretrained VLM backbone
We compare RCORE and the baseline (C2C [li2024c2c]) on Sth-com [li2024c2c] and EK100-com using the InternVideo2-1B [wang2024iv2] backbone, whose 1B-parameter vision encoder is approximately $11\times$ larger than the 87M-parameter vision encoders used in the main paper. In Table 14, RCORE yields consistent improvements over the baseline on Sth-com [li2024c2c] and EK100-com, boosting unseen composition accuracy (+3.4p., +2.1p.), unseen $\Delta_{\text{CG}}$ (+0.8p., +1.1p.), and H.M. (+1.5p., +0.4p.), respectively. These results demonstrate that its powerful representation learning capability is not constrained by the backbone’s parameter size.
0.E.6 Temporal/Static split of Sth-com
@@
We present extended results corresponding to the experiment shown in Figure 5 (c) of the main paper, including all versions of the Temporal/Static split. Inspired by prior work [laura2021temporal, yun2022time], we rigorously evaluate the temporal modeling capabilities by partitioning the Sth-com [li2024c2c] dataset into Temporal and Static splits, as shown in Figure 14. We evaluate both the baseline (C2C [li2024c2c]) and RCORE on the two types of splits, (i) our reconstructed splits and (ii) the splits proposed by Sevilla et al. [laura2021temporal]. In all cases, RCORE exhibits lower performance on shuffled inputs but outperforms the baseline on original inputs. This indicates that RCORE predicts verbs without relying on static cues, leading to robust improvements in verb performance regardless of the split.
In this paragraph, we describe the differences between two Temporal/Static splits. Sevilla et al. [laura2021temporal] originally defined these splits based on cognitive science experiments on Something-Something V2 [goyal2017something] verb classes: a Temporal split (18 verbs) where temporal information matters, and a Static split (18 verbs) where such information is redundant. However, noting that this Static split includes verbs requiring complex temporal reasoning such as ‘Folding’, we reconfigure the splits based on verb semantics. Specifically, we categorized verbs based on this criterion: those requiring temporal information for discrimination are assigned to the Temporal split (62 verbs), while verbs distinguishable by momentary keyframes are assigned to the Static split (80 verbs). The indices of the verbs in the Temporal split are as follows: 0, 1, 5, 6, 14, 25, 26, 29, 30, 32, 33, 34, 35, 38, 39, 40, 41, 42, 43, 44, 45, 46, 58, 59, 60, 61, 65, 66, 67, 68, 72, 73, 74, 75, 76, 78, 79, 85, 86, 90, 91, 92, 98, 100, 103, 109, 122, 123, 131, 134, 136, 137, 139, 140, 145, 146, 147, 148, 149, 150, 153, 154. The indices of the verbs in the Static split are as follows: 2, 3, 8, 9, 10, 11, 12, 13, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 27, 28, 31, 36, 37, 47, 48, 49, 50, 51, 52, 53, 54, 55, 77, 80, 81, 87, 88, 93, 94, 95, 96, 97, 99, 101, 102, 104, 105, 106, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 126, 127, 128, 129, 130, 132, 133, 135, 138, 141, 142, 143, 151, 152, 155, 156, 157, 158, 159, 160. To balance the sample sizes between the Temporal and Static splits, we excluded a few verbs with the fewest samples from the Temporal split.